# Leveraging Label Information for Stealthy Data Stealing in Vertical Federated Learning

### v1

@@ 3.1. VFL protocol
3.2. Threat model
Capability and knowledge. The active client acts as a malicious adversary with the capability to deviate from the standard VFL training protocol. Specifically, this adversary can alter the training task and transmits malicious gradients to other clients. The adversary controls a top model $f_{0}(\cdot)$, and accesses partial data features $\boldsymbol{x}_{j,a},j\in[M]$, along with a bottom model $f_{a}(\cdot)$. Additionally, the adversary possesses an auxiliary dataset, $D_{aux}:=(\boldsymbol{x}_{i},y_{i})_{i=1}^{M_{aux}}$, which shares a similar distribution with the training data but remains distinct from it, i.e., $D_{aux}\cap D_{train}=\emptyset$. Each sample in the auxiliary dataset contains the complete set of features and the corresponding label. This assumption is congruent with previous attacks against SL (Gao and Zhang, 2023; Zhu et al., 2023; Pasquini et al., 2021) and presents a more realistic capability compared to those models that assume direct querying of passive clients’ bottom models. We explicitly do not assume any capability for the adversary to query or directly access the knowledge of the passive clients’ models (see Figure 1).
@@ 3.2. Threat model
4. Challenges and limitations
We point out the main challenges and the limitations of the current attack methods, to achieve effective and stealthy data reconstruction under practical assumptions on VFL settings.
@@ 4. Challenges and limitations
Challenge 2: powerful defenses. Malicious data reconstruction attacks are prone to detection by sophisticated defense mechanisms such as SplitGuard and Gradient Scrutinizer. One major reason is that current malicious attacks, e.g. AGN and FSHA, alter the honest training task and thus affecting the transmitted gradients. While such defenses are highly sensitive to deviations in training behavior, as they typically monitor gradient flows for anomalies that could indicate malicious activities. To make the attack more stealthy, a natural solution is to sum up the honest training task (e.g., classification) loss and the adversarial loss. To evaluate the efficacy of this modification, we employed SplitGuard as a defense mechanism to test against malicious attack strategies AGN and FSHA on MNIST dataset. As depicted in Figure 2, the detection scores for both attacks, even with the modified loss, remained below the threshold of 0.9, yielding successful detection. This outcome suggests that simply combining loss functions is insufficient to penetrate current defenses.
Figure 2. Comparison of Splitguard scores for two malicious attacks, and their variants with modified loss functions.
Limitations: 1) Existing HBC data reconstruction attacks on VFL operates under strong, often unrealistic assumptions about adversary’s capabilities, such as the ability to query the target model or inject data directly into target clients, which are impractical in real-world scenarios; 2) Furthermore, the attacks in SL focus on reconstructing features from a single target client, and do not use any partial features of the data, limiting their applicability in VFL; 3) Malicious attacks like FSHA and AGN completely ignore the label information in their designs, which is the root cause of their detectability, and incorporating the labels into attack designs can potentially further improve the reconstruction performance.
The above challenges and limitations steer us toward a compelling research question: ”Under a practical adversary model, can we develop a malicious data reconstruction attack in VFL that successfully circumvents current defense mechanisms while achieving superior reconstruction performance?” We affirmatively answer this question by developing a novel attack strategy DMAVFL, which addresses all highlighted challenges and limitations.
5. Methodology of DMAVFL
In this section, we introduce our DMAVFL strategy. As illustrated in Figure 3, DMAVFL consists of the following three steps.
  1. (1)
@@ 5. Methodology of DMAVFL
Figure 3. Workflow of DMAVFL. The opaque rectangle indicates that the model is being trained, while the transparent rectangle represents that the model is frozen and only utilized in forward propagation.
5.1. Setup
@@
5.2. Pretraining
Prior to commencing VFL training, the adversary constructs an encoder-decoder structure consisting of the adversary’s bottom model $f_{a}(\cdot)$, the encoder $f_{e}(\cdot)$, and the decoder $f_{d}(\cdot)$. This setup is pretrained on the auxiliary dataset $D_{aux}$. The objective of this pretraining phase is threefold: to train the decoder to reconstruct target features, to enable the adversary’s bottom model to learn the relationship between target and adversary’s partial features, and to allow the encoder to capture the representation of private features for effective reconstruction.
@@ 5.2. Pretraining
(1) $\mathcal{L}_{R}=\frac{1}{|\mathcal{B}_{aux}|}\sum_{i\in\mathcal{B}_{aux}}MSE(\tilde{\boldsymbol{x}}_{i,t},\boldsymbol{x}_{i,t}).$
@@
Algorithm 1 Pretraining.
The pretraining process is detailed in Algorithm 1, where the function Grdient($L,\boldsymbol{\theta}$) computes the gradients of the loss $L$ with respect to parameters $\boldsymbol{\theta}$; and the function Model_update($f(\cdot),\nabla_{\boldsymbol{\theta}}$) represents the operation of updating the model $f(\cdot)$ using gradients $\nabla_{\boldsymbol{\theta}}$.
5.3. Malicious gradient generation
@@
An intuitive solution involves employing a discriminator to distinguish between embeddings from the encoder (labeled as real) and those produced by the target model (labeled as fake). The target model is then trained to deceive the discriminator into recognizing its embeddings as real, thus aligning its distribution with that of the encoder. This strategy focuses on minimizing the Jensen–Shannon divergence (JS) (Fuglede and Topsoe, 2004) between the distributions of the embeddings from the auxiliary and training data, i.e., $\textup{JS}(P_{\boldsymbol{H}_{aux}}\|P_{\boldsymbol{H}_{train}})$, which facilitates the transfer of embeddings distribution.
However, a primary concern is that the discriminator approach completely ignores the label information available at the adversary, which may lead to 1) sub-optimal reconstruction performance; and 2) detection by defenses that can tell the difference between gradients computed using and without using labels. To address these issues, we employ DAC, $D(\cdot)$, defined in (2), which not only differentiates between real and fake embeddings but also distinguishes the corresponding labels.
(2) $\min_{D}\mathbb{E}_{\boldsymbol{h},y\sim P_{\boldsymbol{H}_{aux},Y}}CE(y^{+},D(\boldsymbol{h}))+\mathbb{E}_{\boldsymbol{h},y\sim P_{\boldsymbol{H}_{train},Y}}CE(y^{-},D(\boldsymbol{h})).$
Here, $P_{\boldsymbol{H}_{aux},Y}$ and $P_{\boldsymbol{H}_{train},Y}$ represent the joint distributions of embeddings and labels from the auxiliary and training datasets, respectively. The labels $y^{+}$ and $y^{-}$ denote real and fake labels, respectively, and $\text{CE}(\cdot)$ represents the cross-entropy loss.
@@
(3) $\min_{\boldsymbol{h}}\mathbb{E}_{\boldsymbol{h},y\sim P_{\boldsymbol{H}_{train},Y}}CE(y^{+},D(\boldsymbol{h})).$
@@
Figure 4. t-SNE visualization on Credit dataset with 2 classes.
@@
Figure 5. t-SNE visualization on MNIST dataset with 10 classes.
@@
(4) $\mathcal{L}_{M}=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}CE\left(y_{j}^{+},D(\boldsymbol{h}_{j,p})\right).$
@@
(5) $\mathcal{L}_{D}$ $=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}CE(y_{j}^{-},D(\boldsymbol{h}_{j,p}))$
@@ 5.4. Data reconstruction
6. Experiments
6.1. Experiment setup
@@
Datasets and data processing: We conduct experiments across four representative datasets for VFL tasks, including two tabular datasets (i.e., Credit (Yeh, 2016) and RT_IOT2022 (S. and Nagapadma, 2024)) and two image datasets (i.e., MNIST (Deng, 2012) and CIFAR-10 (Krizhevsky et al., 2009)). Each dataset is partitioned into auxiliary, training, and test sets. For MNIST and CIFAR-10, we use the standard test sets provided. For the Credit dataset, we randomly allocate 30% of the samples to the test set, and for RT_IOT2022, 20% of the samples are selected as the test set. The auxiliary set, chosen randomly from the original training set, comprises 10% of the training set size in all main experiments, ensuring $|D_{aux}|:|D_{train}|=1:10$ and $D_{aux}\cap D_{train}=\emptyset$. Data features are evenly split between passive (target) and active clients, with each holding 50% of the features.
@@
Models: The adversary’s encoder $f_{e}(\cdot)$ is configured identically to the passive clients’ bottom model $f_{p}(\cdot)$ across all datasets. The structure of the adversary’s bottom model $f_{a}(\cdot)$ mirrors that of $f_{p}(\cdot)$. For tabular data, we employ multi-layer perceptron (MLP) models for $f_{e}(\cdot),f_{a}(\cdot)$, $f_{p}(\cdot)$, the decoder $f_{d}(\cdot)$, and the DAC $D(\cdot)$. For image data, convolutional neural networks (CNNs) serve as the foundational architecture, with the CIFAR-10 models featuring a residual block composed of three CNN layers and a skip connection. Detailed dataset information and model structures are provided in the Appendix B.
@@
Metrics: We measure the reconstruction performance using average MSE between the reconstructed features and the original features, denoted as Recon MSE. To analyze the effectiveness of DAC compared to other embedding distribution transfer techniques (e.g., using discriminator or the top model), we measure both the average MSE and the cosine distance between the encoder’s (or shadow model’s) embeddings and the target model’s embeddings, denoted as Emb MSE and Emb Cos, respectively. Before computing the cosine distance, all embeddings are normalized.
@@
Environment: All experiments are conducted on a single machine equipped with four NVIDIA RTX 3090 GPUs. Each experiment is repeated five times, with the average results reported.
@@
Table 1. Results of data reconstruction attacks across four datasets. Attack methods followed by SG or GS represent implementations under SG or GS defenses. DMAVFL and DMAVFL_sync achieve the same performance with and without defenses. Some entries are missing due to 1) some attacks do not use encoders for data reconstruction, and hence do not have embedding MSE and cosine distance; and 2) the GS defense does not work with Credit dataset.
Method Credit RT_IOT2022 MNIST CIFAR-10
@@
Recon
MSE
Emb
MSE
Emb
Cos
Recon
MSE
Emb
MSE
Emb
Cos
Recon
MSE
Emb
MSE
Emb
Cos
GRNA 1.2566 - - 2.2955 - - 0.4209 - - 0.3201 - -
GIA 0.9054 0.2748 0.4949 1.8523 1.3960 0.4077 0.9173 0.0095 0.2535 0.2547 1.1431 0.3210
AGN 1.0934 - - 2.3671 - - 0.0929 - - 0.4549 - -
AGN-SG 1.5348 - - 3.6687 - - 0.5067 - - 0.6535 - -
AGN-GS - - - 4.0843 - - 0.5064 - - 0.4559 - -
PCAT 0.8612 0.5282 0.5486 2.2177 2.3307 0.3817 0.3694 0.0461 0.6016 0.3844 2.4672 0.5544
SDAR 0.5327 0.5451 0.3103 1.6084 2.9362 0.2281 0.0839 0.1638 0.3142 0.1162 1.5072 0.3465
FSHA 0.5152 0.3468 0.1524 1.5150 3.0500 0.4929 0.0156 0.0030 0.0378 0.0317 0.4657 0.1610
FSHA-SG 0.6034 0.5528 0.4043 1.6817 4.3649 0.5839 0.4674 0.0578 0.2016 0.5136 3.5306 0.5516
FSHA-GS - - - 1.8643 2.8935 0.3071 0.5029 0.2879 0.1760 0.1029 1.4048 0.2554
DMAVFL (GS/SG) 0.4389 0.2918 0.1492 1.2789 2.5623 0.0833 0.0152 0.0041 0.0289 0.0296 0.4348 0.1007
DMAVFL_sync (GS/SG) 0.4602 0.3010 0.1529 1.2324 2.1772 0.1025 0.0137 0.0045 0.0388 0.0305 0.4727 0.1121
6.2. Baselines and defenses

### v2

@@ III-A VFL protocol
III-B Threat model
Adversary’s capability and knowledge. In our threat model, the active client is treated as a malicious adversary capable of deviating from the standard VFL training protocol. Specifically, this adversary can alter the training task and transmits malicious gradients to other clients, e.g., alter the classification task to a reconstruction task. As the active client, the adversary controls a top model $f_{0}(\cdot)$, and accesses partial data features $\boldsymbol{x}_{j,a},j\in[M]$, along with a bottom model $f_{a}(\cdot)$ (see Figure 1).
Additionally, the adversary possesses an auxiliary dataset, $D_{aux}:=(\boldsymbol{x}_{i},y_{i})_{i=1}^{M_{aux}}$, which shares a similar distribution with the training data but remains distinct from it, i.e., $D_{aux}\cap D_{train}=\emptyset$. Each sample in the auxiliary dataset contains the complete set of features and the corresponding label. This assumption is congruent with previous attacks against SL [12, 50, 27] and presents a more realistic capability compared to those models that assume direct querying of passive clients’ bottom models. In practice, adversaries often have access to subsets or samples of target data, which are used for validation or other purposes [32]. In some cases, training data may also include a mix of public and private datasets [34, 49], which makes this assumption more practical.
@@ III-B Threat model
IV Challenges
In this section, we point out the main challenges to achieve effective and stealthy data reconstruction attack on VFL. We also review the current attacks’ limitations.
@@ IV Challenges
Challenge 2: powerful detection strategies. Malicious data reconstruction attacks are prone to detection by sophisticated mechanisms such as SplitGuard and Gradient Scrutinizer. One major reason is that current malicious attacks, e.g. AGN and FSHA, alter the honest training task, i.e., replace the honest training loss function with malicious loss function, and thus affecting the transmitted gradients. While such detections are highly sensitive to deviations in training behavior, as they typically monitor gradient flows for anomalies that could indicate malicious activities. To make the attack more stealthy, a natural solution is to integrate the honest training task (e.g., classification) into the malicious task. To evaluate the efficacy of this solution, we employed SplitGuard as a detection mechanism to test against modified malicious attack strategies AGN and FSHA on MNIST dataset, where the training loss is the summation of classifcation loss and the malicious loss. As depicted in Figure 2, the detection scores for both attacks, even with the modified loss, remained below the threshold of 0.9, yielding successful detection. This outcome suggests that simply combining loss functions is insufficient to circumvent current detections.
Figure 2: Comparison of Splitguard scores for two malicious attacks, and their variants with modified loss functions.
Existing HBC data reconstruction attacks on VFL operates under strong, often unrealistic assumptions about adversary’s capabilities, such as the ability to query the target model or inject data directly into target clients, which are impractical in real-world scenarios. Malicious attacks like FSHA and AGN completely ignore the label information in their designs, which is the root cause of their detectability, and incorporating the labels into attack designs can potentially further improve the reconstruction performance.
The above challenges steer us toward a compelling research question: ”Under a practical adversary model, can we develop a malicious data reconstruction attack in VFL that successfully circumvents current detection mechanisms while achieving superior reconstruction performance?” We affirmatively answer this question by developing a novel attack strategy URVFL, which addresses all highlighted challenges.
V Methodology of URVFL
In this section, we introduce our URVFL strategy. As illustrated in Figure 3, URVFL consists of the following three steps.
  1. 1.
@@ V Methodology of URVFL
Figure 3: Workflow of URVFL. The opaque rectangle indicates that the model is being trained, while the transparent rectangle represents that the model is frozen and only utilized in forward propagation.
V-A Setup
@@
V-B Pretraining
Prior to commencing VFL training, the adversary constructs an encoder-decoder structure consisting of the adversary’s bottom model $f_{a}(\cdot)$, the encoder $f_{e}(\cdot)$, and the decoder $f_{d}(\cdot)$. This setup is pretrained on the auxiliary dataset $D_{aux}$. The objective of this pretraining phase is threefold: to train the decoder to reconstruct target features, to enable the adversary’s bottom model to learn the relationship between target and adversary’s partial features, and to allow the encoder to capture the representation of passive clients’ features for effective reconstruction.
@@ V-B Pretraining
$\mathcal{L}_{R}=\frac{1}{|\mathcal{B}_{aux}|}\sum_{i\in\mathcal{B}_{aux}}MSE(\tilde{\boldsymbol{x}}_{i,t},\boldsymbol{x}_{i,t}).$ (1)
@@
Algorithm 1 Pretraining procedure.
The pretraining process is detailed in Algorithm 1, where the function Gradient($L,\boldsymbol{\theta}$) computes the gradients of the loss $L$ with respect to parameters $\boldsymbol{\theta}$; and the function Model_update($f(\cdot),\nabla_{\boldsymbol{\theta}}$) represents the operation of updating the model $f(\cdot)$ using gradients $\nabla_{\boldsymbol{\theta}}$.
V-C Malicious gradient generation
@@
An intuitive solution involves employing a discriminator to distinguish between embeddings from the encoder (labeled as real) and those produced by the target model (labeled as fake). The target model is then trained to deceive the discriminator into recognizing its embeddings as real, thus aligning its distribution with that of the encoder. This strategy focuses on minimizing the Jensen–Shannon divergence (JS) [11] between the distributions of the embeddings from the auxiliary and training data, i.e., $\textup{JS}(P_{\boldsymbol{H}_{aux}}\|P_{\boldsymbol{H}_{train}})$, which facilitates the transfer of embeddings distribution.
However, a primary concern is that the discriminator approach completely ignores the label information available at the adversary, which may lead to 1) sub-optimal reconstruction performance; and 2) detection by methods that can tell the difference between gradients computed using and without using labels. To address these issues, we employ DAC, $D(\cdot)$, defined in (2), which not only differentiates between real and fake embeddings but also distinguishes the corresponding labels.
$\min_{D}\mathbb{E}_{\boldsymbol{h},y\sim P_{\boldsymbol{H}_{aux},Y}}CE(y^{+},D(\boldsymbol{h}))+\mathbb{E}_{\boldsymbol{h},y\sim P_{\boldsymbol{H}_{train},Y}}CE(y^{-},D(\boldsymbol{h})).$ (2)
Here, $P_{\boldsymbol{H}_{aux},Y}$ and $P_{\boldsymbol{H}_{train},Y}$ represent the joint distributions of embeddings and labels from the auxiliary and training datasets, respectively. The labels $y^{+}$ and $y^{-}$ denote real and fake label $y$, respectively, and $\text{CE}(\cdot)$ represents the cross-entropy loss.
@@
$\min_{\boldsymbol{h}}\mathbb{E}_{\boldsymbol{h},y\sim P_{\boldsymbol{H}_{train},Y}}CE(y^{+},D(\boldsymbol{h})).$ (3)
@@
Figure 4: t-SNE visualization on Credit dataset with 2 classes.
@@
Figure 5: t-SNE visualization on MNIST dataset with 10 classes.
@@
$\mathcal{L}_{M}=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}CE\left(y_{j}^{+},D(\boldsymbol{h}_{j,p})\right).$ (4)
@@
$\mathcal{L}_{D}$ $=\frac{1}{|\mathcal{B}_{train}|}\sum_{j\in\mathcal{B}_{train}}CE(y_{j}^{-},D(\boldsymbol{h}_{j,p}))$ (5)
@@ V-D Data reconstruction
VI Experiments
VI-A Experiment setup
@@
Datasets and data processing: We conduct experiments across five representative datasets for VFL tasks, including two tabular datasets (i.e., Credit [48] and RT_IOT2022 [31]) and three image datasets (i.e., MNIST [4], CIFAR-10 [20], and Tiny imagenet [21]). Each dataset is partitioned into auxiliary, training, and test sets. For MNIST and CIFAR-10, we use the standard test sets provided. For the Credit dataset, we randomly allocate 30% of the samples to the test set, and for RT_IOT2022 and Tiny imagenet, 20% of the samples are selected as the test set. The auxiliary set, chosen randomly from the original training set, comprises 10% of the training set size in all main experiments, ensuring $|D_{aux}|:|D_{train}|=1:10$ and $D_{aux}\cap D_{train}=\emptyset$. For image dataset, we normalize the data features to the range [-1,1]. For tabular dataset, most features are continuous or integer. We encode the categorical feature in Credit dataset using sklearn [28] and normalize all features with StandardScaler. Data features are evenly split between passive (target) and active clients, with each holding 50% of the features.
@@
Models: The adversary’s encoder $f_{e}(\cdot)$ is configured identically to the passive clients’ bottom model $f_{p}(\cdot)$ across all datasets. The structure of the adversary’s bottom model $f_{a}(\cdot)$ mirrors that of $f_{p}(\cdot)$. For tabular data, we employ multi-layer perceptron (MLP) models for $f_{e}(\cdot),f_{a}(\cdot)$, $f_{p}(\cdot)$, the decoder $f_{d}(\cdot)$, and the DAC $D(\cdot)$. For image data, convolutional neural networks (CNNs) serve as the foundational architecture, with the CIFAR-10 and Tiny imagenet models featuring a residual block composed of three CNN layers and a skip connection. Detailed dataset information, model structures, and training parameters are provided in the Appendix B.
@@
Metrics: We measure the reconstruction performance using average MSE between the reconstructed features and the original features, denoted as Recon MSE. Besides, to comprehensively measure the reconstruction performance on image datasets, we also leverage PSNR and SSIM as the metrics. To analyze the effectiveness of DAC compared to other embedding distribution transfer techniques (e.g., using discriminator or the top model), we measure both the average MSE and the cosine distance between the encoder’s (or shadow model’s) embeddings and the target model’s embeddings, denoted as Emb MSE and Emb Cos, respectively. Before computing the cosine distance, all embeddings are normalized.
@@
Environment: All experiments are conducted on a single machine equipped with four NVIDIA RTX 4090 GPUs. Each experiment is repeated five times, with the average results reported.
@@
TABLE I: Results of data reconstruction attacks on tabular datasets. Attack methods followed by SG or GS represent implementations under SG or GS detections. URVFL and URVFL_sync achieve the same performance with and without detections. Some entries are missing due to 1) some attacks do not use encoders for data reconstruction, and hence do not have embedding MSE and cosine distance; and 2) the GS defense does not work with Credit dataset.
Method Credit RT_IOT2022
Recon MSE Emb MSE Emb Cos
@@
GRNA $1.1822\pm 0.0541$ $2.2542\pm 0.0325$
GIA $0.9056\pm 0.0002$ $\textbf{0.2738}\pm 0.0008$ $0.4681\pm 0.0215$ $1.8535\pm 0.0012$ $1.7276\pm 0.2356$ $0.4491\pm 0.0371$
AGN $0.9656\pm 0.1196$ $2.2311\pm 0.1256$
AGN-SG $1.4155\pm 0.1734$ $2.3967\pm 0.2628$
AGN-GS - - - $2.5670\pm 0.2909$
PCAT $0.8612\pm 0.0984$ $0.5282\pm 0.0467$ $0.5486\pm 0.0197$ $2.2963\pm 0.4388$ $\textbf{1.5646}\pm 0.2488$ $0.3742\pm 0.0195$
SDAR $0.5327\pm 0.0374$ $0.5451\pm 0.1645$ $0.3103\pm 0.0931$ $1.6084\pm 0.1604$ $2.9362\pm 1.3356$ $0.2281\pm 0.1689$
FSHA $0.5032\pm 0.0382$ $0.3906\pm 0.0622$ $0.2234\pm 0.0564$ $1.5773\pm 0.0507$ $2.2990\pm 0.6506$ $0.4859\pm 0.0984$
FSHA-SG $0.7884\pm 0.1139$ $0.6056\pm 0.0704$ $0.5640\pm 0.0369$ $1.6895\pm 0.0419$ $2.7922\pm 1.0491$ $0.5909\pm 0.0745$
FSHA-GS - - - $1.7442\pm 0.0784$ $2.5768\pm 0.2169$ $0.5534\pm 0.1453$
URVFL (SG/GS) $\textbf{0.4191}\pm 0.0541$ $0.3078\pm 0.0778$ $\textbf{0.1658}\pm 0.0551$ $1.3821\pm 0.0244$ $1.7894\pm 0.7435$ $\textbf{0.0549}\pm 0.0113$
URVFL_sync (SG/GS) $0.4722\pm 0.0228$ $0.3720\pm 0.0462$ $0.2277\pm 0.0441$ $\textbf{1.3277}\pm 0.0665$ $2.3372\pm 0.5489$ $0.0938\pm 0.0340$
TABLE II: Results of data reconstruction on image datasets. Attack methods followed by SG or GS represent implementations under SG or GS detections. URVFL and URVFL_sync achieve the same performance with and without defenses. Some entries are missing since these attacks do not use encoders for data reconstruction, and hence, there is no embedding distance.
Method MNIST CIFAR-10 Tiny imagenet Recon MSE Emb MSE Emb Cos Recon MSE Emb MSE Emb Cos Recon MSE Emb MSE Emb Cos GRNA $0.5533\pm 0.0919$ $0.3287\pm 0.0061$ $0.3869\pm 0.0103$ GIA $0.9125\pm 0.0056$ $0.0179\pm 0.0018$ $0.2772\pm 0.0431$ $0.2550\pm 0.0004$ $1.4011\pm 0.1849$ $0.3957\pm 0.0535$ $0.3234\pm 0.0009$ $1.8109\pm 0.1039$ $0.3936\pm 0.0068$ AGN $0.4801\pm 0.0027$ $0.4413\pm 0.0385$ $0.3786\pm 0.0190$ AGN-SG $0.5131\pm 0.0014$ $0.5154\pm 0.0776$ $1.2398\pm 0.0102$ AGN-GS $0.7073\pm 0.0813$ $0.5901\pm 0.0620$ $1.9577\pm 0.2531$ PCAT $0.2446\pm 0.1312$ $0.0282\pm 0.0016$ $0.3484\pm 0.0299$ $0.3843\pm 0.0123$ $2.4672\pm 0.3024$ $0.5544\pm 0.0023$ $0.3063\pm 0.0249$ $2.7853\pm 0.0521$ $0.5526\pm 0.0025$ SDAR $0.0839\pm 0.0751$ $0.1638\pm 0.1850$ $0.3142\pm 0.1699$ $0.3067\pm 0.0618$ $0.5230\pm 0.6965$ $1.2196\pm 0.4831$ $0.1363\pm 0.0293$ $2.1344\pm 0.1174$ $0.4692\pm 0.0281$ FSHA $0.0536\pm 0.0118$ $0.0168\pm 0.0100$ $0.0909\pm 0.0157$ $0.0317\pm 0.0065$ $0.4657\pm 0.1286$ $0.1610\pm 0.0285$ $0.1025\pm 0.0084$ $1.4897\pm 0.2210$ $0.3056\pm 0.0203$ FSHA-SG $0.5158\pm 0.2511$ $0.0621\pm 0.0095$ $0.2956\pm 0.0691$ $0.1765\pm 0.0390$ $2.1634\pm 0.1310$ $0.4282\pm 0.0480$ $0.3110\pm 0.1202$ $2.8927\pm 0.3759$ $0.4904\pm 0.0519$ FSHA-GS $0.2829\pm 0.1374$ $0.1446\pm 0.0260$ $0.3077\pm 0.0402$ $0.1312\pm 0.0150$ $1.5940\pm 0.0684$ $0.3133\pm 0.0230$ $0.4797\pm 0.2383$ $3.0416\pm 0.4313$ $0.5200\pm 0.0423$ URVFL (SG/GS) $0.0132\pm 0.0027$ $0.0045\pm 0.0009$ $\textbf{0.0287}\pm 0.0029$ $0.0302\pm 0.0011$ $0.5679\pm 0.0834$ $0.1303\pm 0.0131$ $\textbf{0.0699}\pm 0.0055$ $1.4316\pm 0.6673$ $\textbf{0.2139}\pm 0.0195$ URVFL_sync (SG/GS) $\textbf{0.0127}\pm 0.0011$ $\textbf{0.0040}\pm 0.0006$ $0.0341\pm 0.0022$ $\textbf{0.0176}\pm 0.0139$ $\textbf{0.2793}\pm 0.1868$ $\textbf{0.0675}\pm 0.0549$ $0.0704\pm 0.0011$ $\textbf{1.1441}\pm 0.0058$ $0.2434\pm 0.0077$
TABLE III: Results of data reconstruction measured by PSNR and SSIM on image datasets. Attack methods followed by SG or GS represent implementations under SG or GS detections. URVFL and URVFL_sync achieve the same performance with and without defenses. $\uparrow$ denotes that the larger the value, the better the performance.
[…]
VI-B Baselines and detections